# Scaling Methods To Estimate Macroscopic Fundamental Diagrams in Urban Networks with Sparse Sensor Coverage

### v1

@@ 3.1 Statical Scaling
Let’s assume a traffic network ($\Omega$), having links ($i$) of $n$ numbers, with a different hierarchy based on serviceability, defined as link class as $t$. Among all the links, some have loop detectors (LDs) count as $j$. Therefore, the flow in the links of LDs can be presented as $\{q_{j}\;|\;j\in i\}$. The overall network flow $(q_{N})$ and LDs equipped network flow $(q_{eq})$ can be expressed by total travel distance (TTD) as network ‘production’ by all vehicles over the time-space region, shown as follows:
$\bar{q}_{eq}$ $=\frac{\sum_{\forall t}TTD_{eq}}{\sum_{\forall j,t}l_{j,t}\Delta t}$ (1a)
$=\frac{\sum_{\forall j,t}q_{j,t}l_{j,t}}{\sum_{\forall j,t}l_{j,t}}$ (1b)
$\bar{q}_{N}$ $=\frac{\sum_{\forall t}TTD}{\sum_{\forall i,t}l_{i,t}\Delta t}$ (1c)
$=\frac{\sum_{\forall t}TTD_{eq}+\sum_{\forall t}TTD_{neq}}{\sum_{\forall i,t}l_{i,t}\Delta t}$ (1d)
$\sum_{\forall t}TTD_{eq}$ $=\sum_{\forall j,t}q_{j,t}l_{j,t}$ (2)
$\mathbb{E}(q_{j,t}l_{j,t})$ $=\mathbb{E}(q_{j,t})\mathbb{E}(l_{j,t})+Cov(q_{j,t},l_{j,t})$ (3)
$\frac{1}{m}\sum_{\forall t}\sum_{j=1}^{m}q_{j,t}l_{j,t}$ $=\sum_{\forall t}(\frac{1}{m}\sum_{j=1}^{m}q_{j,t})\frac{1}{m}\sum_{\forall t}\sum_{j=1}^{m}l_{j,t}+Cov(q_{j,t},l_{j,t})$ (4)
$\sum_{\forall j,t}q_{j,t}l_{j,t}$ $=\sum_{\forall t}(\bar{q}_{eq,t})\sum_{\forall j,t}l_{j,t}+m\times Cov(q_{j,t},l_{j,t})$ (5)
Similarly, for not equipped network:
$\sum_{\forall t}TTD_{neq}$ $=\sum_{\forall i\notin j,t}q_{i,t}l_{i,t}$ (6)
$\mathbb{E}(q_{i,t}l_{i,t}|\forall i\notin j,t)$ $=\mathbb{E}(q_{i,t}|\forall i\notin j,t)\mathbb{E}(l_{i,t}|\forall i\notin j,t)+Cov(q_{i,t}l_{i,t}|\forall i\notin j,t)$ (7)
$\sum_{\forall t}\sum_{i=1}^{n}q_{i,t}l_{i,t}$ $=\sum_{\forall t}(\bar{q}_{neq,t})\sum_{\forall t}\sum_{i=1}^{n}l_{i,t}+Cov(q_{i,t}l_{i,t}|\forall i\notin j,t)$ (8)
$\sum_{\forall i\notin j,t}q_{i,t}l_{i,t}$ $=\sum_{\forall t}(\bar{q}_{neq,t})\sum_{\forall i\notin j,t}l_{i,t}+n\times Cov(q_{i,t}l_{i,t}|\forall i\notin j,t)$ (9)
@@
Assuming the mean flow and covariance of flow for a unique link hierarchy will be the same. Therefore,
$\bar{q}_{neq,t}=\bar{q}_{eq,t},\hskip 2.84526pt$ $Cov(q_{j,t},l_{j,t})=Cov(q_{i,t}l_{i,t}|\forall i\notin j,t)$ (10)
$\sum_{\forall i\notin j,t}q_{i,t}l_{i,t}$ $=\frac{\sum_{\forall j,t}q_{j,t}l_{j,t}}{\sum_{\forall j,t}l_{j,t}}\sum_{\forall i\notin j,t}l_{i,t}+n\times Cov(q_{j,t},l_{j,t})$ (11)
$=\sum_{\forall j,t}q_{j,t}l_{j,t}\frac{\sum_{\forall i\notin j,t}l_{i,t}}{\sum_{\forall j,t}l_{j,t}}+n\times Cov(q_{j,t},l_{j,t})$ (12)
@@
Thus, the total network flow can be formulated from the equipped network information as follows:
@@
$=\frac{\sum_{\forall j,t}q_{j,t}l_{j,t}+(\sum_{\forall j,t}q_{j,t}l_{j,t}\frac{\sum_{\forall i\notin j,t}l_{i,t}}{\sum_{\forall j,t}l_{j,t}}+n\times Cov(q_{j,t},l_{j,t}))}{\sum_{\forall i,t}l_{i,t}\Delta t}$ (15)
Similarly, we can derive the network average density from the equipped network information. The network average density can be expressed by total travel time (TTT) as the ‘accumulation’ spent by all vehicles in the network over the time-space domain.
@@
$=\frac{\sum_{\forall j,t}k_{j,t}l_{j,t}+(\sum_{\forall j,t}k_{j,t}l_{j,t}\frac{\sum_{\forall i\notin j,t}l_{i,t}}{\sum_{\forall j,t}l_{j,t}}+n\times Cov(k_{j,t},l_{j,t}))}{\sum_{\forall i,t}l_{i,t}\Delta t}$ (18)
The $q_{j}$ and $k_{j}$ are the flow and density in the detector level in a link.
3.2 Geo-spatial Scaling: variogram
Kriging is a well-established geostatistical interpolation method initially developed by Krige [68] and later expanded by Matheron [69]. It has been applied across various disciplines for spatial data interpolation, including missing data imputation in traffic studies. The central component of kriging is the variogram, a function that models the spatial correlation between data points. The variogram provides an essential representation of how spatial dependence changes with distance, guiding Kriging’s ability to estimate unknown values at unobserved locations based on nearby known observations [70].
Variogram models, such as spherical, exponential, and Gaussian models, are typically fitted to the empirical variogram to quantify spatial continuity [71]. These models help define the structure of spatial relationships and enable kriging to make predictions.
@@ 3.2.1 Spatial Dependency and Variogram
Let the traffic flow data for certain links in a network be given as $\{(s_{i},q_{i})\}$, where $s_{i}=(x_{i},y_{i})$ represents the geographical coordinates (latitude and longitude) of the $i$-th known link, and $q_{i}$ is the observed traffic flow at that link. The task is to estimate traffic flow $q(s_{0})$ at unknown locations $s_{0}$ by leveraging spatial interpolation techniques based on kriging and a variogram model.
@@ 3.2.2 Spatial Interpolation: Kriging System Setup
Once the variogram model is fitted, the traffic flow at unmeasured locations can be estimated using Ordinary Kriging. In Ordinary Kriging (OK) for traffic flow estimation, we aim to predict the traffic flow $\hat{q}(s_{0})$ at an unmeasured location $s_{0}$, using a weighted linear combination of the traffic flows $q(s_{i})$ at known locations $s_{i}$, where $i=1,2,\dots,n$. The primary assumption in ordinary kriging is that the mean traffic flow $\bar{q}(s)$ is constant within a neighborhood of the target point $s_{0}$. The following assumptions are considered:
@@ 3.2.3 Ordinary Kriging Estimator:
This completes the estimation procedure for ordinary kriging, where the prediction of traffic flow at a target point is a weighted sum of the flows from nearby measured locations, with weights optimized to minimize prediction error while ensuring the estimator remains unbiased. This ensures that the expected value of the predicted traffic flow is equal to the expected value of the true traffic flow:
@@
The next section describes the empirical data used in this study, focusing on LDD collected under varying levels of sensor deployment. Additionally, it outlines the link hierarchy classification, which forms the basis for the hierarchical scaling methodology proposed in this study.
@@ 4.1 Data Description
The links in the road network were classified in two ways. The first classification follows a three-hierarchy (3-H), where roads are categorized into three groups: Link-1, the most critical roads, and Link-3, the least important major roads. The second classification follows a two-tier hierarchy (2-H), where the network is divided into two types: Link-1, representing the most important roads, and Link-2, which includes the remaining major roads. Notably, Link-1 is identical in both classifications, while Link-2 in the 2-H classification corresponds to Link-2 and Link-3 in the 3-H classification based on their average speeds and flow profiles. Figures 2 and 3 illustrate the comparison of average speeds and flow profiles for the 2-H and 3-H classifications.
@@
The entire network consists of 142 links, spanning different hierarchical levels within the 150 km network. A fully equipped network is defined as one in which all links have at least one LD. In this study, we considered the fully equipped network, which includes 142 LDs distributed across 142 links, as shown in Figures 2 (a, b). To evaluate our methodology, we also created partially equipped networks by randomly removing LDs from various links equally from each hierarchy in both the 2-H and 3-H hierarchical networks. Table 1 provides details on the number of LDs in each link type for the fully equipped (100%) network as well as for networks with 30%, 20%, 10%, and 5% LD coverage. Figure 4 illustrates the distribution of LDs in these partially equipped networks.
In this study, we validated our approach using three different network types: the 2-H network, the 3-H network, and a non-hierarchical network.
@@
(d) 05% LD of Figure 1 network
Figure 4: Partially loop detectors equipped networks (a) 30% LD equipped network, (b) 20% LD equipped network, (c) 10% LD equipped network, and (d) 05 % LD equipped network
@@
3-Hierarchies 2-Hierarchies Non-Hierarchy
Equipped Network Link-1 Link-2 Link-3 Link-1 Link-2 Link-all
@@
Next, section 5 presents the results of applying these methodologies, focusing on the estimation of network demand and MFDs. A comparative analysis of all proposed methods is provided, highlighting their performance under varying levels of sensor deployment.
@@ 5 Results
In this section, we discuss the flow estimation for the whole network from equipped network information. The study proposes two methods: first, statistical scaling of the equipped network and application to the different hierarchical structures of the network. In this study, we demonstrate a two-hierarchical, three-hierarchical, and non-hierarchical network approach to applying the proposed scaling method. Second, the variogram method of spatial scaling, where the empirical variogram calibrates the spherical variogram model using flow distribution in the equipped network and spatially interpolates the flow in the full network. Since the variogram model was developed based on geospatial scaling, we only apply this method to a non-hierarchical network approach.
5.1 Flow Estimation using Geo-spatial methods: variogram
@@
The Variogram model takes input of flow from the equipped network’s links and estimates flow spatially over the full network. Figure 5 demonstrates the flow estimation spatially with respect to different percentages of LD-equipped networks at various times of the day. The color-coded flow plots visually represent traffic flow intensity across the network links, with the gradient indicating varying flow levels, from low (blue) to high (red). Since this study focuses on flow scaling only on the major roads (Link-1, link-2, link-3), we only extracted information along the links in the major roads. The first column of Figure 5, where 30% of the network’s major links are equipped with loop detectors (LD), the Variogram model produces a detailed and accurate estimation of flow across all major and minor links. Congestion hotspots, as indicated by red zones, align closely with expected traffic patterns, validating the model’s capability to capture traffic dynamics effectively under sufficient LD coverage. As the LD percentage is reduced to 20% and 10%, shown in the subsequent columns, the model maintains a good estimation quality of the overall network flow, with significant flow details retained for major road links (Link-1, Link-2, and Link-3). Congestion hotspots and the broader distribution of flow are generally well-captured, indicating the robustness of the Variogram-based estimation even under reduced LD coverage.
@@
When we reduced LD percentage by 20 %, 10% as per Table 1, we could see that the model can estimate the whole network flow along with the major link networks perfectly with the matching of congestion hotspots and other traffic conditions mostly over the different time of the day. As we further decrease the LD percentage to 5%, this method could not able to estimate full network traffic flow because of the sparse information with respect to the variogram radius of estimation. Therefore, this method is best suited for network flow scaling of an equipped network having more than 10 % LD. However, at the 5% LD level (final column), the model’s performance declines. Sparse detector data creates spatial gaps that hinder accurate network flow estimation. This results in less precise representations of flow across the network, with notable discrepancies in congestion hotspots and overall flow intensity. This limitation highlights the importance of maintaining a critical mass of LD coverage for accurate spatial flow estimation.

### v2

@@ 2.1.1 Uniform Scaling or Baseline
Let’s assume a traffic network having links ($i$) of $n$ numbers $(|i|=n)$, with a different hierarchy based on serviceability, defined as link class as $t$. Among all the links, some have loop detectors (LDs) count as $j$. Therefore, the flow in the links of LDs can be presented as $\{q_{j}\;|\;j\in i\}$. The lengths of the links in the network are represented by $l_{i}$. The overall network flow $(\hat{q}_{N})$ can be expressed by total travel distance (TTD) as network ‘production’ by all vehicles over the time-space region, shown as follows in (1):
$\hat{q}_{N}=\frac{\sum_{i=1}^{n}TTD_{i}}{\sum_{i=1}^{n}l_{i}\Delta t}$ (1)
The TTD is the sum of equipped network TTD $(TTD_{j,eq})$ and non-equipped network TTD $(TTD_{i\notin i,neq})$
$TTD=\sum_{i=1}^{n}TTD_{i}=\sum_{\forall j\in\{eq\}}TTD_{j,eq}+\sum_{\forall i\in\{neq\}}TTD_{i,neq}$ (2)
The first term of (2) can be calculated from the LDs.
$\sum_{\forall j\in\{eq\}}TTD_{j,eq}=\sum_{\forall j\in\{eq\}}q_{j}l_{j}$ (3)
The second term of (2) can be directly estimated after the imputation method. Otherwise, we know that the following expectation relation holds for all situations.
$\mathbb{E}(q_{i}l_{i})$ $=\mathbb{E}(q_{i})\mathbb{E}(l_{i})+Cov(q_{i},l_{i})$ (4a)
$\frac{1}{p}\sum_{i\in\{neq\}}(q_{i}l_{i})$ $=\left(\frac{1}{p}\sum_{i\in\{neq\}}q_{i}\right)\frac{1}{p}\sum_{i\in\{neq\}}l_{i}+\times Cov_{neq}(q_{i},l_{i})$ (4b)
$\sum_{i\in\{neq\}}(q_{i}l_{i})$ $=\bar{q}_{neq}\sum_{i\in\{neq\}}l_{i}+p\times Cov_{neq}(q_{i},l_{i})$ (4c)
Here, $\bar{q}_{neq}$ is the mean flow in the non-equipped network, and $p$ is the total number of non-equipped links $(|i\notin j|=p)$. Also, assume $m$ is the total number of equipped links $(|j|=m)$ in the network with mean flow estimated as $\bar{q}_{eq}$ from LDs. If we assume the non-equipped network shares the same average flow and covariance as the equipped, then we can calculate the second term of (2), as (5).
$\sum_{\forall i\in\{neq\}}(q_{i}l_{i})$ $=\bar{q}_{neq}\sum_{\forall i\in\{neq\}}l_{i}+p\times Cov_{neq}(q_{i},l_{i})$ (5a)
$=\left(\frac{1}{m}\sum_{j\in\{eq\}}q_{j}\right)\sum_{\forall i\in\{neq\}}l_{i}+p\times Cov_{eq}(q_{j},l_{j})$ (5b)
$=\bar{q}_{eq}\sum_{\forall i\in\{neq\}}l_{i}+p\times Cov_{eq}(q_{j},l_{j})$ (5c)
Note the covariance term can be disregarded as it was proven negligible in all our data analysis. So, the TTD can be expressed as
$TTD$ $=\bar{q}_{eq}\times(\sum_{\forall i\in\{neq\}}l_{i}+\sum_{\forall j\in\{eq\}}l_{j})$ (6a)
$=\bar{q}_{eq}\times l_{net}$ (6b)
Here $l_{net}$ is the total network length, including equipped and non-equipped links. The assumption, $\bar{q}_{neq}=\bar{q}_{eq}$ forms the foundation of the uniform scaling factor, where all variables are scaled by the factor of network length covered.
Unfortunately, this assumption is often invalid, and equipped links are not a randomly selected subset of the full network. We propose a hierarchical scaling approach that will cluster the equipped network into groups that more likely resemble parts of the non-equipped network.
2.1.2 Hierarchical Network Scaling
The general idea in the hierarchical network scaling method is that we want to approach $\bar{q}_{neq}=\bar{q}_{eq}$ at the cluster level. Similar to (5), we can use the $\bar{q}_{eq,t}$ to estimate the $\bar{q}_{neq,t}$, where $t$ represents link clusters. One option is to use the network hierarchy to define the clusters. Thus, we propose the following scaling method for the hierarchical network approach.
$\sum_{\forall t}TTD_{eq,t}$ $=\sum_{\forall j,t}q_{j,t}l_{j,t}$ (7)
$\mathbb{E}(q_{j,t}l_{j,t})$ $=\mathbb{E}(q_{j,t})\mathbb{E}(l_{j,t})$ (8)
$\frac{1}{m}\sum_{\forall t}\sum_{j=1}^{m}q_{j,t}l_{j,t}$ $=\sum_{\forall t}(\frac{1}{m}\sum_{j=1}^{m}q_{j,t})\frac{1}{m}\sum_{\forall t}\sum_{j=1}^{m}l_{j,t}$ (9)
$\sum_{\forall j,t}q_{j,t}l_{j,t}$ $=\sum_{\forall t}(\bar{q}_{eq,t})\sum_{\forall j,t}l_{j,t}$ (10)
Similarly, for non-equipped networks:
$\sum_{\forall t}TTD_{neq,t}$ $=\sum_{\forall i\in\{neq\}}q_{i,t}l_{i,t}$ (11)
$\mathbb{E}(q_{i,t}l_{i,t}|\forall i\notin j,t)$ $=\mathbb{E}(q_{i,t}|\forall i\notin j,t)\mathbb{E}(l_{i,t}|\forall i\notin j,t)$ (12)
$\sum_{\forall t}\sum_{i=1}^{p}q_{i,t}l_{i,t}$ $=\sum_{\forall t}(\bar{q}_{neq,t})\sum_{\forall t}\sum_{i=1}^{p}l_{i,t}$ (13)
$\sum_{\forall i\notin j,t}q_{i,t}l_{i,t}$ $=\sum_{\forall t}(\bar{q}_{neq,t})\sum_{\forall i\notin j,t}l_{i,t}$ (14)
@@ 2.1.2 Hierarchical Network Scaling
Assuming the mean flow for a non-equipped network is the same as an equipped network for a unique hierarchy, i.e., $(\bar{q}_{neq,t}=\bar{q}_{eq,t})$. Therefore, by applying (10) to (14), we get the following estimation of TTD for a non-equipped network.
$\sum_{\forall i\notin j,t}q_{i,t}l_{i,t}$ $=\frac{\sum_{\forall j,t}q_{j,t}l_{j,t}}{\sum_{\forall j,t}l_{j,t}}\sum_{\forall i\notin j,t}l_{i,t}$ (15)
$=\sum_{\forall j,t}q_{j,t}l_{j,t}\frac{\sum_{\forall i\notin j,t}l_{i,t}}{\sum_{\forall j,t}l_{j,t}}$ (16)
@@
Thus, the total network flow can be estimated from the equipped network information as follows:
@@
$=\frac{\sum_{\forall j,t}q_{j,t}l_{j,t}+(\sum_{\forall j,t}q_{j,t}l_{j,t}\frac{\sum_{\forall i\notin j,t}l_{i,t}}{\sum_{\forall j,t}l_{j,t}})}{\sum_{\forall i,t}l_{i,t}\Delta t}$ (19)
Similarly, we can also derive the network average density from the equipped network. The network average density can be expressed by total travel time (TTT) as the ‘accumulation’ spent by all vehicles in the network over the time-space domain. Similar to the TTD in (2), we can estimate TTT in the equipped network as $TTT_{eq}=\sum_{\forall j}k_{j}l_{j}$. Local densities at LD-level $(k_{j})$ need to be corrected since the LDs suffer from location-biased and systematic errors. This study corrected LD-level density estimation as per the methods mentioned in Maiti (2024). Therefore, the total network density can be formulated from the corrected LD in the equipped network as follows:
@@
$=\frac{\sum_{\forall j,t}k_{j,t}l_{j,t}+(\sum_{\forall j,t}k_{j,t}l_{j,t}\frac{\sum_{\forall i\notin j,t}l_{i,t}}{\sum_{\forall j,t}l_{j,t}})}{\sum_{\forall i,t}l_{i,t}\Delta t}$ (22)
The variables $q_{j,t}$ and $k_{j,t}$ represent flow and density in the equipped network of $t$ link hierarchy, estimated at the detector level on a link.
2.2 Spatial Imputation: Variogram
Instead of scaling observations from the equipped network to represent the non-equipped network, an alternative approach is to impute flow and density values for all links and subsequently calculate the full MFD. To achieve this, we propose using a spatial variogram-based method for the imputation process. Kriging is a well-established geostatistical interpolation method initially developed by Krige (1951) and later expanded by Matheron (1963). It has been applied across various disciplines for spatial data interpolation, including missing data imputation in traffic studies. The central component of kriging is the variogram, a function that models the spatial correlation between data points. The variogram provides an essential representation of how spatial dependence changes with distance, guiding Kriging’s ability to estimate unknown values at unobserved locations based on nearby known observations (Cressie, 1993).
Variogram models, such as spherical, exponential, and Gaussian models, are typically fitted to the empirical variogram to quantify spatial continuity (Oliver, 2014). These models help define the structure of spatial relationships and enable kriging to make predictions.
@@ 2.2.1 Spatial Dependency and Variogram
Let the traffic flow data for certain links in a network be given as $\{(s_{i},q_{i})\}$, where $s_{i}=(x_{i},y_{i})$ represents the geographical coordinates (latitude and longitude) of the $i$-th known link, and $q_{i}$ is the observed traffic flow at that link. The task is to estimate traffic flow $q(s_{0})$ at unknown locations $s_{0}$ by leveraging spatial interpolation techniques based on a variogram-based model.
@@ 2.2.2 Spatial Interpolation: Kriging System Setup
Once the variogram model is fitted, the traffic flow at unmeasured locations can be estimated using Ordinary Kriging. In Ordinary Kriging (OK) for traffic flow estimation, we aim to predict the traffic flow $\hat{q}(s_{0})$ at an unmeasured location $s_{0}$, using a weighted linear combination of the traffic flows $q(s_{i})$ at known locations $s_{i}$, where $i=1,2,\dots,n$. The primary assumption in OK is that the mean traffic flow $\bar{q}(s)$ is constant within a neighborhood of the target point $s_{0}$. The following assumptions are considered:
@@
The expected value of the predicted traffic flow is equal to the expected value of the true traffic flow:
@@ 3 Data
This section describes the empirical data used in this study, focusing on LDD collected under varying levels of sensor deployment. Additionally, it outlines the link hierarchy classification, which forms the basis for the hierarchical scaling methodology proposed in this study.
@@ 3.1 Data Description
The links in the road network were classified in two ways. The first classification follows a three-hierarchy (3-H), where roads are categorized into three groups: Link-1, corresponding to the most critical roads, and Link-3, the least important major roads. The second classification follows a two-tier hierarchy (2-H), where the network is divided into two types: Link-1, representing the most important roads, and Link-2, which includes the remaining major roads. Notably, Link-1 is identical in both classifications, while Link-2 in the 2-H classification corresponds to Link-2 and Link-3 in the 3-H classification based on their average speeds and flow profiles. No additional link classes beyond the 3-H level were identified as exhibiting statistically significant differences in traffic and road characteristics. This link hierarchy is Figures 2 and 3 illustrate the comparison of average speeds and flow profiles for the 2-H and 3-H classifications. We can see in Figure 3 that the average speed and flow in each group are significantly different, so using a common flow average from the equipped network to estimate the average flow on the non-equipped network leads to strong bias.
@@
The study focuses on a road network comprising 2,456 links, spanning various hierarchical levels within a 150 km area. A fully equipped network is defined as one where every link is equipped with at least one LD. However, only 142 of these links currently have LDs installed. Therefore, the LDs percentage in the network is approximately 5.78%. To address this limitation, a partially equipped network was constructed, utilizing the data from these 142 links to estimate MFDs for validation purposes. For comparison, the fully equipped network was conceptualized as including the 142 LDs distributed across their respective links, as illustrated in Figures 2 (a, b). To evaluate our methodology, we also created partially equipped networks by randomly removing LDs from various links equally from each hierarchy in both the 2-H and 3-H hierarchical networks. Table 1 provides details on the number of LDs in each link type for the fully equipped (100%) network as well as for networks with 30%, 20%, 10%, and 5% LD coverage. Figure 4 illustrates the distribution of LDs in these partially equipped networks.
In this study, the proposed approach was validated using two distinct network types: the 2-H network, the 3-H network, and a baseline non-hierarchical network for comparison.
@@
(d) 5% LD of Figure 1 network
Figure 4: Partially loop detectors equipped networks (a) 30% LD equipped network, (b) 20% LD equipped network, (c) 10% LD equipped network, and (d) 5 % LD equipped network
@@
Equipped 3-Hierarchies 2-Hierarchies Non-Hierarchy
Network Link-1 Link-2 Link-3 Link-1 Link-2 Link-all
@@ 4 Results
The study proposes two methods: first, hierarchical scaling with application to the 2-H and 3-H network approach. Second, the spatial data imputation with the variogram method to spatially interpolate the network variables in the full network. In the end, we compared the estimated network variables and MFDs with the baseline non-hierarchical network approach.
4.1 Flow estimation using spatial imputation methods: variogram
@@
The Variogram model takes input of flow from the equipped network’s links and estimates flow spatially over the full network. Figure 5 demonstrates the flow estimation spatially with respect to different percentages of LD-equipped networks at various times of the day. The color-coded flow plots visually represent traffic flow intensity across the network links, with the gradient indicating varying flow levels, from low (blue) to high (red). Since this study focuses on flow scaling only on the major roads (Link-1, link-2, link-3), we only extracted information along the links in the major roads. In the first column of Figure 5, where 30% of the network’s major links are equipped with loop detectors (LD), the Variogram model produces a detailed and accurate estimation of flow across all major and minor links. Congestion hotspots, as indicated by red zones, align closely with expected traffic patterns, validating the model’s capability to capture traffic dynamics effectively under sufficient LD coverage. As the LD percentage is reduced to 20% and 10%, shown in the subsequent columns, the model maintains a good estimation quality of the overall network flow, with significant flow details retained for major road links (Link-1, Link-2, and Link-3). Congestion hotspots and the broader distribution of flow are generally well-captured, indicating the robustness of the Variogram-based estimation even under reduced LD coverage.
@@
As we further decrease the LD percentage to 5%, this method was not able to estimate full network traffic flow because of the sparse information with respect to the variogram radius of estimation. Therefore, this method is best suited for network flow scaling of an equipped network having more than 10 % LD. However, at the 5% LD level (final column), the model’s performance declines. Sparse detector data creates spatial gaps that hinder accurate network flow estimation. This results in less precise representations of flow across the network, with notable discrepancies in congestion hotspots and overall flow intensity. This limitation highlights the importance of maintaining a critical mass of LD coverage for accurate spatial flow estimation. Our goal is not to determine the flow on links but to perform imputation in a way that ensures accurate aggregation when calculating the network’s average flow.